# Aragog: Just-in-Time Model Routing for Scalable Serving of Agentic Workflows

### v1

@@ 6 Additional Related Work
LLM serving optimizations The rapid deployment of large language models has driven significant innovations in serving optimizations. Some focus on application level compression that craft lightweight variants for efficient inference [awq, gptq, llm-pruner, speculative-decoding]. Some employ compiler techniques to optimize execution for low latency [mirage, mlc-llm]. Other works focus on system perspectives like request scheduling [orca, llmschedulinglearning], KV cache memory management [vllm, jenga], parallelism [alpaserve, nanoflow, magescale-infer], efficient kernels [flashattention, flashinfer], and resource allocation [llumnix, distserve, prism]. Aragog operates as an orchestration layer between LLM serving engines and the agentic workflow layer. As such, these optimizations are orthogonal to Aragog and can be incorporated into the underlying serving engines.
Agent serving optimizations The rise of LLM-powered agents introduces new challenges for LLM serving, as agent execution involves an input-dependent and varying number of interceptions (e.g., tool calls) and LLM calls. Recent research has focused on addressing these agent-specific inefficiencies [autellix, infercept, pdm, tempo, asyncfunccall]. For instance, InferCept [infercept] reduces GPU resource waste arising from frequent LLM interceptions, while Autellix [autellix] mitigates head-of-line blocking caused by the variable number of LLM calls per agent by prioritizing requests based on cumulative service time. These techniques are complementary to Aragog, which orchestrates execution at the workflow level (across a graph of agents) and can leverage agent-serving optimizations within each agent within the graph.
Agentic workflow serving optimizations Agentic workflows have emerged as a more reliable and debuggable approach than fully autonomous agents. These workflows orchestrate multiple agents as a graph, introducing dependencies across agents. Recent systems exploit graph-level knowledge for scheduling and resource management [parrot, ayo, kairos, batchquery]. For instance, Parrot [parrot] enables graph-aware batching and KV cache sharing, while Ayo [ayo] rearranges the execution graph for efficient execution. These systems optimize scheduling and placement for a fixed configuration; Aragog complements them by dynamically adapting the configuration at runtime. Orthogonally, other work automates the design and search of workflow graph structures [aflow, flow]. Aragog is agnostic to the input workflow’s structure and can efficiently serve these automatically generated workflows.
@@ 7 Conclusion
We present Aragog, an efficient agentic workflow serving system that progressively adapt request’s configuration to runtime dynamics. Aragog’s core insight is to decouple accuracy and cost based configuration selection. The expensive accuracy-based configuration prediction happens once and before execution, while the lightweight cost-based configuration selection runs frequently throughout execution. Aragog makes runtime reconfiguration efficient with novel acceleration strategies. Across diverse workloads, Aragog achieves 50.0–217.0% higher serving capacity than baselines while maintaining the accuracy of the most expensive configuration. Aragog delivers lower serving latency under different request loads. Under peak load, it delivers 32.5–78.9% and 46.8–81.5% lower median and P95 latency, respectively.

### v2

@@ 6 Additional Related Work
LLM serving optimizations. The rapid deployment of large language models has driven significant innovations in serving optimizations. Some focus on application level compression that craft lightweight variants for efficient inference [awq, gptq, llm-pruner, speculative-decoding]. Some employ compiler techniques to optimize execution for low latency [mirage, mlc-llm]. Other works focus on system perspectives like request scheduling [orca, llmschedulinglearning], KV cache memory management [vllm, jenga], parallelism [alpaserve, nanoflow, magescale-infer], efficient kernels [flashattention, flashinfer], and resource allocation [llumnix, distserve, prism]. Aragog operates as an orchestration layer between LLM serving engines and the agentic workflow layer. As such, these optimizations are orthogonal to Aragog and can be incorporated into the underlying serving engines.
Agent serving optimizations. The rise of LLM-powered agents introduces new challenges for LLM serving, as agent execution involves an input-dependent and varying number of interceptions (e.g., tool calls) and LLM calls. Recent research has focused on addressing these agent-specific inefficiencies [autellix, infercept, pdm, tempo, asyncfunccall]. For instance, InferCept [infercept] reduces GPU resource waste arising from frequent LLM interceptions, while Autellix [autellix] mitigates head-of-line blocking caused by the variable number of LLM calls per agent by prioritizing requests based on cumulative service time. These techniques are complementary to Aragog, which orchestrates execution at the workflow level (across a graph of agents) and can leverage agent-serving optimizations within each agent within the graph.
Agentic workflow serving optimizations. Agentic workflows orchestrate multiple agents as a graph, introducing dependencies across agents. Recent systems exploit graph-level knowledge for scheduling and resource management [parrot, ayo, kairos, batchquery]. For instance, Parrot [parrot] introduces graph-aware batching and KV cache sharing, while Ayo [ayo] rearranges the execution graph for efficient execution and scheduling. These systems optimize for a fixed configuration; Aragog complements them by dynamically adapting configurations at runtime. Other work automates workflow graph design [aflow, flow]; Aragog is agnostic to graph structure and can serve automatically-generated workflows.
@@ 7 Conclusion
We present Aragog, an efficient agentic workflow serving system that progressively adapts a request’s configuration to runtime dynamics. Aragog’s core insight is to decouple accuracy and cost based configuration selection. The expensive accuracy-based configuration prediction happens once and before execution, while the lightweight cost-based configuration selection runs frequently throughout execution. Aragog makes runtime reconfiguration efficient with novel acceleration strategies. Across diverse agentic workflows and model families, Aragog achieves 42.8–217.0% higher serving capacity than baselines while maintaining the accuracy of the most expensive configuration. Aragog delivers lower serving latency under different request loads. Under peak load, it delivers 32.5–86.1% and 46.8–89.0% lower median and P95 latency, respectively.